OpenStudy (kirbykirby):

Find the sum \[\sum_{k=1}^{x}\frac{1}{2^k}\]

OpenStudy (kirbykirby):

The answer says it's \[1-\frac{1}{2^x}\] , but I'm not sure how they get it. Is it not the formula for a finite geometric series? a = 1, r = 1/2 so \[1 (\frac{1-\frac{1}{2^x}}{1-\frac{1}{2}})= 2-2^{1-x}?\]

OpenStudy (anonymous):

\[ \frac 1 2 + \frac 1 {2^2} + \cdots + \frac 1 {2^x} =\\ \frac 1 2 \left ( 1+ \frac 1 {2^1} + \cdots + \frac 1 {2^{x-1}} \right)= \frac 1 2\frac{1-\frac{1}{2^x}}{ \left(1-\frac{1}{2}\right)}=1-2^{-x} \] The first term is \( \frac 1 2\) and not 1.

OpenStudy (kirbykirby):

oohh I see ok. My brain is dead right now I've been up for way too long :( Thanks for the help :)

OpenStudy (anonymous):

yw